Circular No. 6800

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

GRB 971227 A. J. Castro-Tirado and J. Gorosabel, Laboratorio de Astrofísica Es-

pacial y Física Fundamental, Madrid; J. Greiner, Astrophysikalisches Institut, Potsdam; M. R. Zapatero-Osorio, Instituto de Astrofisica de Canarias, Tenerife; and E. Costa, Istituto di Astrofisica Spaziale, Frascati (on behalf of the BeppoSAX team) report: “Following the detection of GRB 971227 by BeppoSAX and BATSE (IAUC 6796, 6798), R-band images were obtained by Y. Aguilar and R. Kohley (Sternwarte der Universität, Bonn) with the 2.2-m CAHA telescope (+ CAFOS) at the German-Spanish Calar Alto Observatory. Only the image taken on Dec. 27.91 UT includes the entire error box of 1SAX J1257.3+5924, presumably the GRB 971227 x-ray afterglow (IAUC 6797). An object with R = 19.5 (preliminary magnitude) at the edge of the 16′ field-of-view, and inside the 1SAX J1257.3+5924 error box, is seen on the Dec. 27 image at α = 12^{h}57^{m}10^{s}.6, δ = +59^{o}24′43′′ (equinox 2000.0; uncertainty ± 3′′), but it is not present on images taken on Dec. 30 at Loiano (see below), to a limiting magnitude of R = 20.5. An identification chart can be obtained via anonymous ftp at laeff.esa.es under /pub/users/ajct/grb971227. Deeper observations are encouraged in order to confirm whether this is the optical counterpart of GRB 971227.”

C. Bartolini, A. Guarnieri, M. Lolli, A. Piccioni, and F. Zavatti, Università di Bologna, communicate: “The object proposed by Castro-Tirado et al. (above) is undetectable on our R frames on Dec. 30.16, which were taken at the 1.5-m telescope of the Bologna Astronomical Observatory in Loiano. The object was therefore fainter than R = 20.5 (mag limit of the images, preliminary value). Data reduction is in progress. Deeper imaging is highly recommended.”

SUPERNOVA 1997ei IN NGC 3963 K. Ayani, Bisei Astronomical Observatory (BAO); and H. Yamaoka,

Kyushu University, report: “A spectrum of SN 1997ei was obtained with the BAO 1.01-m telescope on Dec. 28.8 UT, from which we confirm the classification (type Ia) reported on IAUC 6796. However, a Si II feature near 615 nm is rather shallow compared with typical (bright) type-Ia supernovae. A Si II feature around 580 nm is also visible. These features suggest that SN 1997ei is a somewhat peculiar type-Ia object, probably fast-declining and sub-luminous (Nugent et al. 1995, Ap.J. 455, L147).”

1997 December 30  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green